OpenStudy (anonymous):

in an isolated town of 5000 inhabitants,160 people have a disease at the beginning of the week and 1200 have it at the end of the week. How long does it take for 80% of the population to become infected? Using differential equations.

OpenStudy (anonymous):

What , like P= P_0 e^kt, like that?

OpenStudy (anonymous):

f(t)=160e^(kt) let t be measured in days when t=7, f(x)=1200 1200=160e^(7k) 1200/160=e^(7k) 7.5=e^(7k) ln(7.5)=7k ln(7.5)/7= k =.2878 Now we have the equation f(x)=160e^(.2878t) So, when 80% are infected, 4000 are infected, so solve for t 4000=160e^(.2878t) 25=e^(.2878t) ln(25)=.2878t ln(25)/.2878=t=11.184

OpenStudy (anonymous):

Basically, you solve for k (which is a constant) given the initial conditions of a certain value of f(x) at a given time. then solve for t when f(x)=4000

OpenStudy (anonymous):

oh ok finally an answer

OpenStudy (anonymous):

cause the way its written in txtbook it uses partial fractions then integration

OpenStudy (anonymous):

That's the textbook probably showing you how you get the general form of the equation.

OpenStudy (anonymous):

ohhhhh ok thanks

OpenStudy (dumbcow):

im not sure how to set it up using differential equations, however you can set up a function that models this situation \[x(0) = 160, x(7) = 1200, x(\infty) = 5000\] \[x(t) = \frac{5000}{1+30.25e^{kt}}\] setting x(7) = 1200 and solving for k \[k = -.3224\] 80% of population is 4000 set x(t)=4000 and solve for t \[t = 14.875\] So it takes about 15 days for 80% to become infected